OpenStudy (cherryfrost27):

Simones age is 4/5 of her sisters age. What fraction of the total age of both girls is her sisters age?

OpenStudy (cherryfrost27):

@Benji

OpenStudy (benji):

Okay, if Simone's sister is 25 (I chose 25 because it is divisible by 5), how old is Simone if she is 4/5 her sister's age?

OpenStudy (cherryfrost27):

so would i divide 25 by 4/5?

OpenStudy (cherryfrost27):

wait but isnt 20 divisible by 5 too?

OpenStudy (benji):

20 is divisible by 5 too, yes, but I chose 25 because I cheated and it's also divisible by 4, haha. If 25/5 = 5, than what is 4/5 of 25, or 4*5?

OpenStudy (cherryfrost27):

im not sure

OpenStudy (benji):

If her sister is 25, and Simone is 4/5 of her age, than she would be 20, because 25/5=5 and 4*5=20

OpenStudy (benji):

So Simone is 20 and her sister is 25, what is the total if their ages are put together?

OpenStudy (cherryfrost27):

45

OpenStudy (benji):

Yes, so what is the fraction of her sisters age over the sum of both their ages?

OpenStudy (cherryfrost27):

25/45

OpenStudy (benji):

Exactly, that can be simplified a little though, both are divisible by 5.

OpenStudy (cherryfrost27):

ok 5/9?

OpenStudy (benji):

Yep, So Simone is 4/5 of her sisters age, her sister is 5/9 of their total age.

OpenStudy (cherryfrost27):

OOOHHH okay thanks so much again!

OpenStudy (benji):

yep

OpenStudy (cherryfrost27):

:)